OpenStudy (anonymous):

Give the value of the maximum or minimum of the function f(x) = x2 - 1.

OpenStudy (zarkon):

\[0\le x^2\] \[-1\le x^2-1\]

OpenStudy (campbell_st):

easiest way is to use x = -b/2a in your question a = 1, b = 0 and c = 1 x = 0/2 x = 0... a line of symmetry at x = 0.... (the min or max in on that line) sub into the original equation gives y = -1 the parabola is concave up so the minimum is at (0, -1)

OpenStudy (anonymous):

awesome thank you, guys